Factor completely
100-c^6

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

remember that a^2-b^2=(a+b)*(a-b)
100=a^2 so a=10
b^2=c^6=(c^3)^2
(10-c^3)*(10+c^3)

Answer by solver91311(24713)





Factor the difference of two squares:



Bearing in mind that the cube root of 10 cubed is 10, factor the sum and difference of two cubes:



You can also apply the quadratic formula to the two trinomials and get the following result:



Which is the end because you started with a 6th degree polynomial and ended with 6 first degree binomial factors.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism